OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x)=x^2+81x+7 from x = 4 to x = 7.

OpenStudy (amoodarya):

\[average \space rate =\frac{f(b)-f(a)}{b-a}\] in this case a=4 b=7 so find \[\frac{f(7)-f(4)}{7-4}\]